2. x + 6x + 18x sin 2t, x(0) = -1, x0) = 1.
Answer #1

Sun) x+6x + 18z = sinet , * = -1, (0:1 let d at - Sinat So associated i- (02+60 +18) Homogeneous equation (52 +6D +18) x =P. I. = 340 *o ? - vingt-o (w21)oz? PT.= 1 = > Il 14 sin 2t - 12 60327 340 Tsinat - 6caat? 170 and General solution is 3 = Co